### v1

# Phase diagram of the interacting Haldane model with spin-dependent sublattice potentials

Abstract
Using the exact-diagonalization (ED) and mean-field (MF) approaches, we study the ground-state phase diagram of the interacting Haldane model on the honeycomb lattice, taking into consideration of the spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$. Here $\alpha=\text{A}$,B and $\sigma=\uparrow$$\downarrow$ indicate the sublattice and spin components, respectively. If we set $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$), the system favors a spin ordered state. On the other hand, introducing the nearest-neighbor Coulomb interaction $V$ can drive the system to be charge ordered. Due to their competition, we find that in both ED and MF approaches, an exotic state with Chern number $C=1$ survives from two locally ordered phases and a topologically ordered phase with $C=2$. In ED method, other properties, including the fidelity metric, the excitation gap and the structure factors, are also used to determine the critical points. In MF method with large enough lattice size, we define the local order parameters and band gap to characterize the phase transitions. The interacting Haldane model and the spin-dependent lattice potential can be realized in an ultracold atom gas, which may serve as a way to detect this intriguing state.
I Introduction
Different from the traditional framework of the Landau-Ginzburg theory that relies on locally defined order parameters resulting from broken symmetries, topological phases have been identified and characterized based on their global, nonlocal propertiesHasan and Kane (2010); Qi and Zhang (2011). Over the past few years, the categorization of topologically ordered states in non-interacting systems has been completed, in terms of different symmetriesChiu et al. (2016); Zhang et al. (2019); Kruthoff et al. (2017); Vergniory et al. (2019); Tang et al. (2019); Schnyder et al. (2008).
On the other hand, interacting topological insulators, which are known as systems encompassing the interplay of topological properties and electronic correlations, have been extensively studied in recent years Rachel (2018); Hohenadler and Assaad (2013). Correlation effects are expected to generate exotic states in the presence of topologically nontrivial conditions, such as the antiferromagnetic topological state in the Bernevig-Hughes-Zhang model with the on-site Hubbard interactionMiyakoshi and Ohta (2013); Yoshida et al. (2013) and the topologically non-trivial phase with $C=1$ ($C$ denotes the Chern number) in the spinful Haldane-Hubbard model on honeycomb latticeHe et al. (2011); Zhu et al. (2014); Wu et al. (2015); Vanhala et al. (2016); Tupitsyn and Prokof’ev (2019); Mertz et al. (2019); Yuan et al. (2023). The origin of this $C=1$ phase is attributed to a spontaneous SU(2) symmetry breaking with one of the spin components in the Hall state and the other one in a localized state. Similar studies include the antiferromagnetic Chern insulator in Kane-Mele-Hubbard model Jiang et al. (2018) and $C=1$ phase in interacting topological models on the square lattice Wang and Qi (2019); Ebrahimkhas et al. (2021). Other attempts like introducing the double exchange processes to the Haldane Hamiltonian can also break the spin symmetry Tran and Tran (2022), while interplays between topology, on-site and nearest-neighbor interactions do not exhibit such phenomenon Wang and Qi (2019); Shao et al. (2021). In experiment, observing these states are still challenging within the field of material science, and thus the trapped cold atoms in controlled optical lattice may provide an alternative and promising way for achieving this purpose. So far, the experimental realization of the topological Haldane model has been reportedJotzu et al. (2014) and the quantum simulations of strongly correlated systems (including the Fermi-Hubbard model) in ultracold Fermi gases can be reviewed in Ref. Hofstetter and Qin (2018); Esslinger (2010); Tarruell and Sanchez-Palencia (2018).
In this paper, inspired by the studies of spin-dependent optical latticeJotzu et al. (2015); Förster et al. (2009); Mandel et al. (2003), we propose an interacting Haldane model on honeycomb lattice with spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$, where $\sigma=\uparrow,\downarrow$ and $\alpha=$A,B represents the spin and sublattice components, respectively. If one set $\Delta_{\uparrow,\text{A}}=+\Delta$, $\Delta_{\uparrow,\text{B}}$$=-\Delta$, $\Delta_{\downarrow,\text{A}}=-\Delta$ and $\Delta_{\downarrow,\text{B}}$$=+\Delta$, the system favors a spin-density-wave (SDW) state. Meanwhile, we introduce the nearest-neighbor Coulomb interaction $V$ to drive the system into a charge-density-wave (CDW) state. Due to their competitions, an intermediate state can be expected and we thus study its ground-state phase diagram. We find that except for the topologically non-trivial phase with Chern number $C=2$ and two topologically trivial phases (SDW and CDW) with $C=0$, an newly generated topological phase with $C=1$ can be observed in both exact-diagonalization (ED) and mean-field (MF) method. Our finding provide a new version of competitions among topology, electronic correlation and lattice potential to realize the exotic state.
The presentation is organized as follows: we introduce the model, methods and relevant quantities in Sec. II. Sections III presents our results based on the ED and MF approaches. Lastly, a conclusion is given in Sec. IV.
II Model and measurements
The Hamiltonian of the interacting Haldane model with spin-dependent lattice potentials can be writen as
$\hat{H}=\hat{H}_{k}+\hat{H}_{l},$ (1)
where the kinetic part
[…]
and the local part
$\hat{H}_{l}$ $=$ $U\sum_{i}\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}+V\sum_{\langle i,j\rangle,\sigma,\sigma^{\prime}}\hat{n}_{i,\sigma}\hat{n}_{i,\sigma^{\prime}}$ (3)
$+$ $\sum_{i,\sigma}\Delta_{\sigma,\alpha}\hat{n}_{i,\sigma}.$
In Eq. (2), $\hat{c}^{\dagger}_{i,\sigma}$ ($\hat{c}_{i,\sigma}$) is the creation (annihilation) operator for an electron at site $i$ with spin $\sigma=\uparrow{\rm or}\downarrow$. $t_{1}$ ($t_{2}$) is the nearest-neighbor (next-nearest-neighbor) hopping constant and the Haldane phase $\phi_{i,j}=\phi$ ($-\phi$) in the clockwise (anticlockwise) loop is introduced to the next-nearest-neighbor hopping terms. In Eq. (3), $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ is the number operator; $U$ and $V$ are the on-site and nearest-neighbor Coulomb interaction strength, respectively. $\Delta_{\sigma,\alpha}$ is the spin-dependent lattice potential: for spin $\sigma=\uparrow$, $\Delta_{\sigma,A}=+\Delta$ and $\Delta_{\sigma,B}=-\Delta$; for spin $\sigma=\downarrow$, $\Delta_{\sigma,A}=-\Delta$ and $\Delta_{\sigma,B}=+\Delta$.
In what follows, the model in Eq. (1) are named as the extended Haldane-Hubbard model with spin-dependent lattice potentials. Throughout the paper, we set $t_{1}=1$, $t_{2}=0.2$, $\phi=\pi/2$ and focus on the ground-state phase diagram of this model at half-filling.
II.1 Exact diagonalization in real space
The topological invariant is one of the most improtant properties to characterize the topological phase transitions, which can be quantified by the Chern number in our model. Given the twisted boundary conditions (TBCs) Poilblanc (1991), it can be evaluated by Niu et al. (1985),
$C=\int\frac{d\phi_{x}d\phi_{y}}{2\pi{\rm i}}\left(\langle\partial_{\phi_{x}}\Psi|\partial_{\phi_{y}}\Psi\rangle-\langle{\partial_{\phi_{y}}\Psi|\partial_{\phi_{x}}\Psi\rangle}\right),$ (4)
with ${|\Psi\rangle}$ being the ground-state wave function. Here $\phi_{x}$ and $\phi_{y}$ are the twisted phases along two directions. To avoid the integration of the wave function ${|\Psi\rangle}$ with respect to the continuous variables, we instead use a discretized version Fukui et al. (2005); Varney et al. (2011); Zhang et al. (2013) with intervals $\Delta\phi_{x}=2\pi/N_{x}$ and $\Delta\phi_{y}=2\pi/N_{y}$. In what follows, ($N_{x}$, $N_{y}$)=(20, 20) is adopted to calculate the Chern number.
Other properties used to characterize the critical behavior of quantum phase transition include the ground-state fidelity metric $g$, which is defined as Zanardi and Paunković (2006); Campos Venuti and Zanardi (2007); Zanardi et al. (2007)
$g(x,\delta x)\equiv\frac{2}{N}\frac{1-|\langle\Psi(x)|\Psi(x+\delta x)\rangle|}{(\delta x)^{2}},$ (5)
where x represents the parameters V or $\Delta$, and N is the lattice size. $|\Psi(x)\rangle$ [$|\Psi(x+\delta x)\rangle$] is the ground state of $\hat{H}(x)$ [$\hat{H}(x+\delta x)$] and we set $\delta x=10^{-3}$. In addition, the SDW and CDW structure factors can be used to characterize the spin and charge ordered insulators, respectively. Their definitions in a staggered fashion can be written as
$S_{\text{SDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}-\hat{n}_{j,\downarrow})\rangle,$
$S_{\text{CDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}+\hat{n}_{j,\downarrow})\rangle,$ (6)
where $\eta=0$ ($\eta=1$) if sites $i$ and $j$ are in the same (different) sublattice.
II.2 Mean-field method in momentum space
We employ a variational mean-field method in momentum space to contrast the results obtained via the ED method, which has been reported in studying the extended Haldane-Hubbard model without lattice potentials Shao et al. (2021). In our case, by introducing the operators $a_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{A}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$ and $b_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{B}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$, the Hamiltonian (1) can be expressed as
$\hat{H}=\hat{H}_{\text{0}}+\hat{H}_{\text{I}},$ (7)
where
$\hat{H}_{\text{0}}=\sum_{\textbf{k},\sigma}\left(m_{+,\sigma}(\textbf{k})a_{\textbf{k},\sigma}^{{\dagger}}a_{\textbf{k},\sigma}+m_{-,\sigma}(\textbf{k})b_{\textbf{k},\sigma}^{{\dagger}}b_{\textbf{k},\sigma}\right.$
$\phantom{=\;\;}\left.-t_{1}g(\textbf{k})a_{\textbf{k},\sigma}^{{\dagger}}b_{\textbf{k},\sigma}-t_{1}g^{*}(\textbf{k})b_{\textbf{k},\sigma}^{{\dagger}}a_{\textbf{k},\sigma}\right),$ (8)
and
$\hat{H}_{\text{I}}$ $=\frac{U}{N}\sum_{\textbf{k},\textbf{k'},\textbf{q}}c_{\textbf{k+q},\uparrow}^{{\dagger}}c_{\textbf{k},\uparrow}c_{\textbf{k}^{\prime}-\textbf{q},\downarrow}^{{\dagger}}c_{\textbf{k}^{\prime},\downarrow}$
$+\frac{V}{N}\sum_{\sigma,\sigma^{\prime}}\sum_{\textbf{k},\textbf{k'},\textbf{q}}g(\textbf{q})a_{\textbf{k+q},\sigma}^{{\dagger}}a_{\textbf{k},\sigma}b_{\textbf{k}^{\prime}-\textbf{q},\sigma^{\prime}}^{{\dagger}}b_{\textbf{k}^{\prime},\sigma^{\prime}}.$ (9)
Here $g(\textbf{k})=1+e^{-\mathrm{i}\textbf{k}\cdot\textbf{a}_{1}}+e^{-\mathrm{i}\textbf{k}\cdot\textbf{a}_{2}}$, and we set
$m_{+,\uparrow}(\mathbf{k})$ $=+\Delta+m_{+}(\mathbf{k})$
$m_{+,\downarrow}(\mathbf{k})$ $=-\Delta+m_{+}(\mathbf{k})$
$m_{-,\uparrow}(\mathbf{k})$ $=-\Delta+m_{-}(\mathbf{k})$
$m_{-,\downarrow}(\mathbf{k})$ $=+\Delta+m_{-}(\mathbf{k})$ (10)
with $m_{\pm}(\mathbf{k})=-2t_{2}[\cos({\bf k}\cdot{\bf a}_{1}\mp\phi)+\cos({\bf k}\cdot{\bf a}_{2}\pm\phi)+\cos({\bf k}\cdot({\bf a}_{1}-{\bf a}_{2})\pm\phi)]$.
By decoupling the four-fermion terms in Eq.(9), the mean-field Hamiltonian can be written as
$\hat{H}_{\rm MF}=\hat{H}_{0}+\sum_{\textbf{k}}\psi_{\textbf{k}}^{{\dagger}}\left(\begin{array}[]{cccc}\varepsilon_{\uparrow}^{a}&\xi_{\uparrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow\downarrow}^{a}&\xi_{\uparrow\downarrow}(\textbf{k})\\ \xi^{*}_{\uparrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow}^{b}&\xi^{*}_{\downarrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow\downarrow}^{b}\\ (\varepsilon_{\uparrow\downarrow}^{a})^{*}&\xi_{\downarrow\uparrow}(\textbf{k})&\varepsilon_{\downarrow}^{a}&\xi_{\downarrow\downarrow}(\textbf{k})\\ \xi^{*}_{\uparrow\downarrow}(\textbf{k})&(\varepsilon_{\uparrow\downarrow}^{b})^{*}&\xi^{*}_{\downarrow\downarrow}(\textbf{k})&\varepsilon_{\downarrow}^{b}\end{array}\right)\psi_{\textbf{k}},$
where $\psi^{\dagger}_{\textbf{k}}=[a^{\dagger}_{\textbf{k},\uparrow},b^{\dagger}_{\textbf{k},\uparrow},a^{\dagger}_{\textbf{k},\downarrow},b^{\dagger}_{\textbf{k},\downarrow}]$ is the basis for each lattice momentum ${\bf k}$ and
$\xi_{\sigma\sigma^{\prime}}(\textbf{k})$ $=-\frac{V}{N}\sum_{\textbf{q}}g(\textbf{k}-\textbf{q})\langle b_{\textbf{q},\sigma^{\prime}}^{{\dagger}}a_{\textbf{q},\sigma}\rangle_{\rm MF},$
$\varepsilon_{\sigma}^{a}$ $=Un_{-\sigma}^{a}+3V\sum_{\sigma^{\prime}}n_{\sigma^{\prime}}^{b},$
$\varepsilon_{\sigma}^{b}$ $=Un_{-\sigma}^{b}+3V\sum_{\sigma^{\prime}}n_{\sigma^{\prime}}^{a},$
$\varepsilon_{\uparrow\downarrow}^{a}$ $=-\frac{U}{N}\sum_{\textbf{q}}\langle a_{\textbf{q},\downarrow}^{{\dagger}}a_{\textbf{q},\uparrow}\rangle_{\rm MF},$
$\varepsilon_{\uparrow\downarrow}^{b}$ $=-\frac{U}{N}\sum_{\textbf{q}}\langle b_{\textbf{q},\downarrow}^{{\dagger}}b_{\textbf{q},\uparrow}\rangle_{\rm MF},$ (11)
with densities $n_{\sigma}^{a}=\frac{1}{N}\sum_{\textbf{q}}\langle a_{\textbf{q},\sigma}^{{\dagger}}a_{\textbf{q},\sigma}\rangle_{\rm MF}$ and $n_{\sigma}^{b}=\frac{1}{N}\sum_{\textbf{q}}\langle b_{\textbf{q},\sigma}^{{\dagger}}b_{\textbf{q},\sigma}\rangle_{\rm MF}$. The above mean-field equations can be solved self-consistently by making use of the variational mean-field approach. Once the free energy has converged, the SDW and CDW order parameters can be obtained by
$\mathcal{O}_{\text{SDW}}$ $=$ $\left|\frac{1}{2}\left(\langle\vec{S}_{A}\rangle_{\rm MF}-\langle\vec{S}_{B}\rangle_{\rm MF}\right)\right|,$
$\mathcal{O}_{\text{CDW}}$ $=$ $\left|\left(n^{A}_{\uparrow}+n^{A}_{\downarrow})-(n^{B}_{\uparrow}+n^{B}_{\downarrow}\right)\right|.$ (12)
Here $\vec{S}_{i}=\frac{1}{2}\sum_{\alpha\beta}c_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}c_{i\beta}$, and $\vec{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the vector of spin-1/2 Pauli matrices. Meanwhile, we use the discrete formulation in its multiband (non-Abelian) version to compute the Chern number Fukui et al. (2005).
III Results and analysis
III.1 Results of the Exact Diagonalization
In our ED calculations, we choose the $12$A cluster whose reciprocal lattice contains the $\Gamma$, K, ${}^{\prime}$ points, and one pair of M points, see more details in Ref. Shao et al. (2021). It has been shown that containing the high-symmetry K and ${}^{\prime}$ points are critical to study the quantum phase transition in interacting Haldane systems Shao et al. (2021); Varney et al. (2010, 2011). Meanwhile, we employ the periodic boundary condition and make use of translational symmetries to reduce the Hilbert space size. Therefore, discussions with regard to the excited states below are restricted to the momentum subspace where the ground state belongs to.
The ($V$, $\Delta$) phase diagrams are shown in Figs. 1(a), 1(b), 1(c) and 1(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. The phase transition points (black circles) in Fig. 1 are identified by the positions of the peaks of fidelity metric $g$. Notice that in Fig. 1(a) with $U=0.0$, the results of Chern number matches the phase diagram obtained from the fidelity metric very well, where we use the blue, green and red squares to represent the Chern number $C=2$, $C=1$ and $C=0$, respectively. We can observe from Fig. 1 (a) that the Chern insulator (CI) phase with $C=2$ is in a region with small $V$ and $\Delta$, the spontaneous symmetry breaking SDW or CDW phases with $C=0$ dominates the system when $\Delta$ or $V$ is large enough. Interestingly, an intermediate phase with $C=1$ can be observed surrounded by three expected phases. It is different from the phase diagram in Ref. Vanhala et al. (2016), where the $C=1$ phase is sandwiched between a band insulator and Mott insulator. Notice that their charge ordered phase (band insulator) is governed by the ionic potential $\Delta_{\rm AB}$ which is spin-independent, and their spin ordered phase (Mott insulator) is governed by the on-site Hubbard interaction $U$. While in our case, the charge ordered phase (CDW) is governed by the nearest-neighbor interaction $V$ and the spin ordered phase (SDW) is governed by the spin-dependent lattice potential $\Delta_{\sigma,\alpha}$. We speculate the different competition fashions result in the different locations of $C=1$ phase. In addition, we observe that the $C=1$ phase in Fig. 1(a) can exist with very small $V$, while in small $\Delta$ region ($\Delta\leq 0.5$) it vanishes. This is completely opposite to our results of MF method in Sec. III.2, where the $C=1$ phase can exist in small $\Delta$ region rather than small $V$ region, see Fig. 3(a). We suggest that this is an open issue and expect further studies on it.
Figure 1: Phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of fidelity metric $g$ with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$. The blue, green and red squares indicate results of Chern number $C=2$, $C=1$ and $C=0$, respectively. The black dashed lines in (a) and (d) denote the parameters we choose to show more details below.
When $U=1.0$ in Fig. 1(b), however, results of Chern number do not match the phase diagram obtained from the fidelity metric. Similar discrepancies between Chern number and fidelity metric have also been discussed in Ref. Yuan et al. (2023) and they found that increasing $(N_{x},N_{y})$ from (6, 6) to (20, 20) can effectively reduce such discrepancies. Remember that $N_{x}$ and $N_{y}$ come from the discretized method to calculate Chern number, see the relevant discussions in Sec. II. While here we already use $(N_{x}=20,N_{y}=20)$ to calculate the Chern number, and results of $C=1$ (green points) can not even be observed in Figs. 1(c) and 1(d) with $U=2.0$ and $U=3.0$, respectively. We further use $(N_{x}=60,N_{y}=60)$ to check several points in Fig. 1(d) but with no finding of $C=1$. On the other hand, the area of the intermediate phase identified by the fidelity metric decreases but does not vanishes as $U$ increases from $0.0$ to $3.0$. (We must have a statement here that in our ED discussions, the $C=1$ phase is equivalent to the intermediate phase when $U=0.0$, and they are not equivalent when $U$ is finite). The MF results in Sec. III.2 also show the existence of $C=1$ in the case of $U=3.0$. We thus suggest the issue of whether the $C=1$ phase can exist with $U$ increased to 3.0 may still be open. This issue is worth of further study by other methods like Monte Carlo and density matrix renormalization group because in materials $U$ is generally larger than $V$. Back to Fig. 1(a), the situation with $U=0.0$ and finite $V$ may still be realized in ultracold atom systems due to their advances in parameter simulations.
Figure 2: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $V=0.4$ on the left panels and $\Delta=1.4$ on the right panels. The on-site interaction $U=0$ and the parameters are corresponding to the balck dashed lines in Fig. 1(a).
To detail the critical behaviors of the phase diagram, we select two balck dashed lines with $V=0.4$ and $\Delta=1.4$ in Fig. 1(a) to calculate the relevant properties. The first four lowest-lying energy levels $E_{\alpha}$ ($\alpha=0,1,2,3$ and $\alpha=0$ represents the ground state) are obtained by employing the Arnoldi Lehoucq et al. (1997) method, see Figs. 2(a) and 2(e). In Fig. 2(a) with $V=0.4$, the system crosses the $C=2$, $C=1$ and SDW phases as $\Delta$ increases from $0.6$ to $1.4$, and two level crossings between the ground state and one excited state can be observed at $\Delta\approx 0.9$ and $\Delta\approx 1.2$. In Fig. 2(e) with $\Delta=1.4$, the system crosses the SDW, $C=1$ and CDW phases as $V$ increases from $0.0$ to $2.0$, and two level crossings locates at $V\approx 0.75$ and $V\approx 1.25$. It may be more intuitive to exhibit the excitation gaps $\Delta_{\rm ex}^{1}=E_{1}-E_{0}$ and $\Delta_{\rm ex}^{2}=E_{2}-E_{0}$, see Figs. 2(b) and 2(f). We can find that the excitation gaps exhibit local minimum values at the critical points where level crossings occur. It should be pointed out that in Fig. 2(f), $\Delta_{\rm ex}^{1}$ becomes very small when $V>1.25$, which is due to the almost degenerate ground state and first excited state in CDW phase. Another feature to characterize the phase transitions is the change of the CDW or SDW structure factor ($S_{\text{CDW}}$ or $S_{\text{SDW}}$) in Figs. 2(c) and 2(g). Values of $S_{\text{CDW}}$ ($S_{\text{SDW}}$) in $C=1$ phase is larger than those in $C=2$ phase but smaller than those in CDW (SDW) phase, which indicating that $C=1$ is an intermediate state as a result of the interplay between topology, electronic correlations and lattice potentials. In addition, sharp peaks of the ground-state fidelity metrics $g$ in Figs. 2(d) and 2(h) can be also observed and used to characterize the critical points. Finally, notice that the left and right panels in Fig. 2 are corresponding to the results of $V=0.4$ and $\Delta=1.4$, respectively.
For $U=3.0$ in Fig. 1(d), similar calculations and discussions along the black dashed lines of $\Delta=0.4$ and $\Delta=0.6$ are detailed in Appendix A.
III.2 Results of the Mean-Field Method
To contrast the exact results on small lattice above, we now report the outcomes of the MF method. By implementing a variational mean-field approximation, the lattice size can be enlarged significantly. We use the $180\times 180$ lattice to calculate the band gap and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), while a $30\times 30$ lattice is adopted for the calculations of Chern number. Even though their lattice size are different, critical points identified by the Chern number and other properties are consistent, as will be discussed in Fig. 4. Phase diagrams with regard to the parameters $V$ and $\Delta$ are shown in Figs. 3(a), 3(b), 3(c) and 3(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. Compared to the phase diagrams of ED in Fig. 1, similar features can be observed: SDW and CDW dominate the system for large $\Delta$ and $V$, respectively, leaving the $C=2$ phase in small ($V$, $\Delta$) region; the $C=1$ phase are surrounded by the above three phases. Differently, the increasing interaction $U$ suppresses but does not eliminate the $C=1$ state.
Figure 3: Mean-field phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of Chern number $C$ and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$.
Taking the $U=0.0$ case in Fig. 3(a) as an example, we show in Figs. 4(a) and 4(b) the CDW and SDW order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), respectively, as a function of $V$ and $\Delta$. We find that $\mathcal{O}_{\text{CDW}}$ vanishes in SDW and $C=2$ phases, while the $C=1$ phase can be regarded as a zone of transition for $\mathcal{O}_{\text{CDW}}$ decreases from $2$ to 00. Notice that the phase boundary between SDW and CDW can only be identified from the order parameters, and other phase boundaries are characterized by the results of Chern number, see Fig. 4(d). Even though there are some defect points in Fig. 4(d) which may be due to the severe quantum fluctuation in the vicinity of the critical points, the phase boundaries can be readily identified. For $\mathcal{O}_{\text{SDW}}$ in Fig. 4(b), the $C=1$ phase is still a transition region with finite values but smaller that those in SDW phase. These features are similar to the structure factors of ED results shown in Figs. 2(c) and 2(g). To contrast the excitation gaps in ED results, we show the band gap defined as $\Delta(\textbf{k})$=min [$E_{2}(\textbf{k}^{\prime})-E_{1}(\textbf{k}^{\prime})$] in Fig. 4(c). The rapid drop or closing of the gap size can be observed at the critical values except for the phase boundary between SDW and CDW, where only discontinuity of gap size appears instead. This can be connected with the general picture that the change of a topological invariant is always accompanied by a single-particle gap closing. For the phase transition from SDW to CDW side, the Chern number does not change ($C=0$). That may be a reason to explain the discontinuity of gap size instead of gap closing, and remember that they are all gapped phases with different gap sizes.
Figure 4: Contour plots of (a) the CDW order parameter, (b) the SDW order parameter, (c) the band gap and (d) the Chern number, as a function of $V$ and $\Delta$. The mean-field approach is adopted for the calculations of the model (1) with $U=0.0$.
IV Summary and Discussion
To summarize, we studied the interacting spinful Haldane model at half-filling on the honeycomb lattice with spin-dependent sublattice potentials. By employing the exact-diagonalization (ED) and mean-field (MF) method, we obtained the similar ground-state phase diagrams with regard to the nearest-neighbor interaction $V\in[0,3]$ and sublattice potential difference $\Delta\in[0,3]$: the CDW and SDW phases are in the large $V$ and $\Delta$ region, respectively; the $C=2$ phase dominates the system when both $V$ and $\Delta$ are small enough; a intermediate phase with $C=1$ are surrounded by the above three phases. Except for the change of topological invariant, i.e., the Chern number, other features like the closure of excitation gap, change of structure factors and peak of the fidelity metric in ED results are also observed. As a contrast, the mean-field Chern number, band gap and local order parameters are also studied, which makes the existence of such $C=1$ phase more evident.
On the other hand, the interaction $U$ suppresses the $C=1$ phase but leaving an open question of whether it can exist with $U$ increased to $3.0$. In ED, $C=1$ can not be found when $U=3.0$ even though an intermediate phase can still be identified by results of the fidelity metric. In MF, $C=1$ phase preserves for $U=3.0$ even though its area is very small. Further studies on this interesting issue are thus expected.
Acknowledgements.
C. S. acknowledges support from the National Natural Science Foundation of China (NSFC; Grant No. 12104229) and the Fundamental Research Funds for the Central Universities (Grant No. 30922010803). H.-G. L. acknowledges support from NSFC (Grants No. 11834005 and No. 12247101), and the National Key Research and Development Program of China (Grant No. 2022YFA1402704).
Appendix A details of some properties with $U=3$
Figure 5: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $\Delta=0.4$ on the left panels and $\Delta=0.6$ on the right panels. The on-site interaction $U=3$ and the parameters are corresponding to the balck dashed lines in Fig. 1(d).
In the case of $U=3.0$, we select $\Delta=0.4$ (left panel) and $\Delta=0.6$ (right panel) to present the relevant properties in Fig. 5. Remember that we do not observe the results of $C=1$ in Fig. 1(d) with $U=3.0$. The line of $\Delta=0.4$ ($\Delta=0.6$) crosses the $C=2$ phase (SDW phase), the intermediate phase and the CDW phase as $V$ increases from $1.0$ to $2.0$ [see black dashed lines in Fig. 1(d)]. From the ground-state fidelity metrics $g$ in Figs. 5(d) and 5(h), we can find the sharp peaks between the intermediate and the CDW phases. Accompanied with the shape peaks, level crossings between the ground state and second excited state [Figs. 5(a) and 5(e)], minimum of the second excitation gap $\Delta_{\text{ex}}^{(2)}$ [Figs. 5(b) and 5(f)] and sudden changes of structure factors [Figs. 5(c) and 5(g)] can be observed. Notice that in CDW phase the ground state and first excited state are almost degenerate. On the other hand, only “hump”s can be observed at the phase boundaries between the $C=2$ (or SDW) and the intermediate phase. Near the positions of the “hump”s, evolutions of the low-energy levels are more complicated and changes of structure factors are more smooth. Such features have also been discussed in Refs. Shao et al. (2021); Yuan et al. (2023), where its reason is attributed to the finite-size effect and it can be eliminated by using different clusters or twisted boundary conditions.